Please help me solve the equation:
11-5(3-x)=-2(x+3)+6x

Answer by mananth(16946)

11-5(3-x)=-2(x+3)+6x
11-15+5x=-2x+6+6x
+5x+2x-6x=-11+15+6
x=10